Two beams, A and B, made with the same material are to be subjected to a three-point bending test. Beam A has a dimension of 1.25 in x 2.5 in x 0.5 in while Beam B has a dimension of 1.5 in x 1.5 in x 0.75 in. Compare the respective fracture loads that will cause the beams to fail.
1
Expert's answer
2020-12-07T03:46:18-0500

For beam A we have dimension as L"_A=" 2.5 inch, B"_A=1.5 , H_A=0.5 inch"

for flexural modulus we know that

"E_F= \frac{L^3 m}{bd^3}= \frac{2.5^3 \times m}{1.5\times 0.5^3}=41.67 m"

Similary, for the beam B which has dimension given as "1.5\times 1.5\times 0.75"

"E_F= \frac{L^3 m}{bd^3}= \frac{1.5^3 \times m}{1.5\times 0.75^3}= 4 m"

So here we are getting different flexural modulus for beam A and beam B

The flexural modulus of A is more than that of B